\begin{thebibliography}{10}

\bibitem{abboudtalg2018}
Amir Abboud, Arturs Backurs, Thomas~Dueholm Hansen, Virginia Vassilevska~Williams, and Or~Zamir.
\newblock Subtree isomorphism revisited.
\newblock {\em ACM Trans. Algorithms}, 14(3), jun 2018.
\newblock \href {https://doi.org/10.1145/3093239} {\path{doi:10.1145/3093239}}.

\bibitem{abboudsoda2016}
Amir Abboud, Arturs Backurs, Thomas~Dueholm Hansen, Virginia~Vassilevska Williams, and Or~Zamir.
\newblock {\em Subtree Isomorphism Revisited}, pages 1256--1271.
\newblock URL: \url{https://epubs.siam.org/doi/abs/10.1137/1.9781611974331.ch88}, \href {https://arxiv.org/abs/https://epubs.siam.org/doi/pdf/10.1137/1.9781611974331.ch88} {\path{arXiv:https://epubs.siam.org/doi/pdf/10.1137/1.9781611974331.ch88}}, \href {https://doi.org/10.1137/1.9781611974331.ch88} {\path{doi:10.1137/1.9781611974331.ch88}}.

\bibitem{abboudfocs2015}
Amir Abboud, Arturs Backurs, and Virginia~Vassilevska Williams.
\newblock Tight hardness results for lcs and other sequence similarity measures.
\newblock In {\em 2015 IEEE 56th Annual Symposium on Foundations of Computer Science}, pages 59--78, 2015.
\newblock \href {https://doi.org/10.1109/FOCS.2015.14} {\path{doi:10.1109/FOCS.2015.14}}.

\bibitem{abboudstocs2016}
Amir Abboud, Thomas~Dueholm Hansen, Virginia~Vassilevska Williams, and Ryan Williams.
\newblock Simulating branching programs with edit distance and friends: or: a polylog shaved is a lower bound made.
\newblock In {\em Proceedings of the Forty-Eighth Annual ACM Symposium on Theory of Computing}, STOC '16, page 375–388, New York, NY, USA, 2016. Association for Computing Machinery.
\newblock \href {https://doi.org/10.1145/2897518.2897653} {\path{doi:10.1145/2897518.2897653}}.

\bibitem{amirfocs2014}
Amir Abboud and Virginia~Vassilevska Williams.
\newblock Popular conjectures imply strong lower bounds for dynamic problems.
\newblock In {\em 2014 IEEE 55th Annual Symposium on Foundations of Computer Science}, pages 434--443, 2014.
\newblock \href {https://doi.org/10.1109/FOCS.2014.53} {\path{doi:10.1109/FOCS.2014.53}}.

\bibitem{abboudicalp2014}
Amir Abboud, Virginia~Vassilevska Williams, and Oren Weimann.
\newblock Consequences of faster alignment of sequences.
\newblock In Javier Esparza, Pierre Fraigniaud, Thore Husfeldt, and Elias Koutsoupias, editors, {\em Automata, Languages, and Programming}, pages 39--51, Berlin, Heidelberg, 2014. Springer Berlin Heidelberg.

\bibitem{abeliukalgorithms2013}
Andrés Abeliuk, Rodrigo Cánovas, and Gonzalo Navarro.
\newblock Practical compressed suffix trees.
\newblock {\em Algorithms}, 6(2):319--351, 2013.
\newblock URL: \url{https://www.mdpi.com/1999-4893/6/2/319}, \href {https://doi.org/10.3390/a6020319} {\path{doi:10.3390/a6020319}}.

\bibitem{adjerohbook2008}
Donald Adjeroh, Timothy Bell, and Amar Mukherjee.
\newblock {\em The Burrows-Wheeler Transform: Data Compression, Suffix Arrays, and Pattern Matching}.
\newblock Springer Publishing Company, Incorporated, 1 edition, 2008.

\bibitem{tasuya1993}
Tatsuya Akutsu.
\newblock A linear time pattern matching algorithm between a string and a tree.
\newblock In Alberto Apostolico, Maxime Crochemore, Zvi Galil, and Udi Manber, editors, {\em Combinatorial Pattern Matching}, pages 1--10, Berlin, Heidelberg, 1993. Springer Berlin Heidelberg.

\bibitem{amir2000}
Amihood Amir, Moshe Lewenstein, and Noa Lewenstein.
\newblock Pattern matching in hypertext.
\newblock {\em Journal of Algorithms}, 35(1):82--99, 2000.
\newblock URL: \url{https://www.sciencedirect.com/science/article/pii/S0196677499910635}, \href {https://doi.org/10.1006/jagm.1999.1063} {\path{doi:10.1006/jagm.1999.1063}}.

\bibitem{apostolicocaw1985}
Alberto Apostolico.
\newblock The myriad virtues of subword trees.
\newblock In Alberto Apostolico and Zvi Galil, editors, {\em Combinatorial Algorithms on Words}, pages 85--96, Berlin, Heidelberg, 1985. Springer Berlin Heidelberg.

\bibitem{backursstoc2015}
Arturs Backurs and Piotr Indyk.
\newblock Edit distance cannot be computed in strongly subquadratic time (unless seth is false).
\newblock In {\em Proceedings of the Forty-Seventh Annual ACM Symposium on Theory of Computing}, STOC '15, page 51–58, New York, NY, USA, 2015. Association for Computing Machinery.
\newblock \href {https://doi.org/10.1145/2746539.2746612} {\path{doi:10.1145/2746539.2746612}}.

\bibitem{backursfocs2016}
Arturs Backurs and Piotr Indyk.
\newblock Which regular expression patterns are hard to match?
\newblock In {\em 2016 IEEE 57th Annual Symposium on Foundations of Computer Science (FOCS)}, pages 457--466, 2016.
\newblock \href {https://doi.org/10.1109/FOCS.2016.56} {\path{doi:10.1109/FOCS.2016.56}}.

\bibitem{backursscicomp2018}
Arturs Backurs and Piotr Indyk.
\newblock Edit distance cannot be computed in strongly subquadratic time (unless seth is false).
\newblock {\em SIAM Journal on Computing}, 47(3):1087--1097, 2018.
\newblock \href {https://arxiv.org/abs/https://doi.org/10.1137/15M1053128} {\path{arXiv:https://doi.org/10.1137/15M1053128}}, \href {https://doi.org/10.1137/15M1053128} {\path{doi:10.1137/15M1053128}}.

\bibitem{bakerstoc1993}
Brenda~S. Baker.
\newblock A theory of parameterized pattern matching: algorithms and applications.
\newblock In {\em Proceedings of the Twenty-Fifth Annual ACM Symposium on Theory of Computing}, STOC '93, page 71–80, New York, NY, USA, 1993. Association for Computing Machinery.
\newblock \href {https://doi.org/10.1145/167088.167115} {\path{doi:10.1145/167088.167115}}.

\bibitem{beckerspire2023}
Ruben Becker, Davide Cenzato, Sung-Hwan Kim, Bojana Kodric, Alberto Policriti, and Nicola Prezza.
\newblock Optimal wheeler language recognition.
\newblock In Franco~Maria Nardini, Nadia Pisanti, and Rossano Venturini, editors, {\em String Processing and Information Retrieval}, pages 62--74, Cham, 2023. Springer Nature Switzerland.

\bibitem{billecpm2022}
Philip Bille, Inge~Li G{\o}rtz, Shay Mozes, Teresa~Anna Steiner, and Oren Weimann.
\newblock {The Fine-Grained Complexity of Episode Matching}.
\newblock In Hideo Bannai and Jan Holub, editors, {\em 33rd Annual Symposium on Combinatorial Pattern Matching (CPM 2022)}, volume 223 of {\em Leibniz International Proceedings in Informatics (LIPIcs)}, pages 4:1--4:12, Dagstuhl, Germany, 2022. Schloss Dagstuhl -- Leibniz-Zentrum f{\"u}r Informatik.
\newblock URL: \url{https://drops.dagstuhl.de/entities/document/10.4230/LIPIcs.CPM.2022.4}, \href {https://doi.org/10.4230/LIPIcs.CPM.2022.4} {\path{doi:10.4230/LIPIcs.CPM.2022.4}}.

\bibitem{bringmannfocs2014}
Karl Bringmann.
\newblock Why walking the dog takes time: Frechet distance has no strongly subquadratic algorithms unless seth fails.
\newblock In {\em Proceedings of the 2014 IEEE 55th Annual Symposium on Foundations of Computer Science}, FOCS '14, page 661–670, USA, 2014. IEEE Computer Society.
\newblock \href {https://doi.org/10.1109/FOCS.2014.76} {\path{doi:10.1109/FOCS.2014.76}}.

\bibitem{bringmannfocs2015}
Karl Bringmann and Marvin Künnemann.
\newblock Quadratic conditional lower bounds for string problems and dynamic time warping.
\newblock In {\em 2015 IEEE 56th Annual Symposium on Foundations of Computer Science}, pages 79--97, 2015.
\newblock \href {https://doi.org/10.1109/FOCS.2015.15} {\path{doi:10.1109/FOCS.2015.15}}.

\bibitem{burrows1994}
M.~Burrows and D.~J. Wheeler.
\newblock A block-sorting lossless data compression algorithm.
\newblock Technical report, 1994.

\bibitem{cliffordipl2007}
Peter Clifford and Raphaël Clifford.
\newblock Simple deterministic wildcard matching.
\newblock {\em Information Processing Letters}, 101(2):53--54, 2007.
\newblock URL: \url{https://www.sciencedirect.com/science/article/pii/S002001900600250X}, \href {https://doi.org/10.1016/j.ipl.2006.08.002} {\path{doi:10.1016/j.ipl.2006.08.002}}.

\bibitem{colestoc2002}
Richard Cole and Ramesh Hariharan.
\newblock Verifying candidate matches in sparse and wildcard matching.
\newblock In {\em Proceedings of the Thiry-Fourth Annual ACM Symposium on Theory of Computing}, STOC '02, page 592–601, New York, NY, USA, 2002. Association for Computing Machinery.
\newblock \href {https://doi.org/10.1145/509907.509992} {\path{doi:10.1145/509907.509992}}.

\bibitem{colescicomp2003}
Richard Cole and Ramesh Hariharan.
\newblock Tree pattern matching to subset matching in linear time.
\newblock {\em SIAM Journal on Computing}, 32(4):1056--1066, 2003.
\newblock \href {https://arxiv.org/abs/https://doi.org/10.1137/S0097539700382704} {\path{arXiv:https://doi.org/10.1137/S0097539700382704}}, \href {https://doi.org/10.1137/S0097539700382704} {\path{doi:10.1137/S0097539700382704}}.

\bibitem{cotumacciojacm2023}
Nicola Cotumaccio, Giovanna D’Agostino, Alberto Policriti, and Nicola Prezza.
\newblock Co-lexicographically ordering automata and regular languages - part i.
\newblock {\em J. ACM}, 70(4), aug 2023.
\newblock \href {https://doi.org/10.1145/3607471} {\path{doi:10.1145/3607471}}.

\bibitem{cotumacciosoda2021}
Nicola Cotumaccio and Nicola Prezza.
\newblock {\em On Indexing and Compressing Finite Automata}, pages 2585--2599.
\newblock URL: \url{https://epubs.siam.org/doi/abs/10.1137/1.9781611976465.153}, \href {https://arxiv.org/abs/https://epubs.siam.org/doi/pdf/10.1137/1.9781611976465.153} {\path{arXiv:https://epubs.siam.org/doi/pdf/10.1137/1.9781611976465.153}}, \href {https://doi.org/10.1137/1.9781611976465.153} {\path{doi:10.1137/1.9781611976465.153}}.

\bibitem{crochemorebook2021}
Maxime Crochemore, Thierry Lecroq, and Wojciech Rytter.
\newblock {\em 125 Problems in Text Algorithms: with Solutions}.
\newblock Cambridge University Press, 2021.

\bibitem{dubinerfocs1990}
M.~Dubiner, Z.~Galil, and E.~Magen.
\newblock Faster tree pattern matching.
\newblock In {\em Proceedings [1990] 31st Annual Symposium on Foundations of Computer Science}, pages 145--150 vol.1, 1990.
\newblock \href {https://doi.org/10.1109/FSCS.1990.89533} {\path{doi:10.1109/FSCS.1990.89533}}.

\bibitem{dubinerjacm1994}
Moshe Dubiner, Zvi Galil, and Edith Magen.
\newblock Faster tree pattern matching.
\newblock {\em J. ACM}, 41(2):205–213, mar 1994.
\newblock \href {https://doi.org/10.1145/174652.174653} {\path{doi:10.1145/174652.174653}}.

\bibitem{durajipec2017}
Lech Duraj, Marvin K\"{u}nnemann, and Adam Polak.
\newblock {Tight Conditional Lower Bounds for Longest Common Increasing Subsequence}.
\newblock In Daniel Lokshtanov and Naomi Nishimura, editors, {\em 12th International Symposium on Parameterized and Exact Computation (IPEC 2017)}, volume~89 of {\em Leibniz International Proceedings in Informatics (LIPIcs)}, pages 15:1--15:13, Dagstuhl, Germany, 2018. Schloss Dagstuhl -- Leibniz-Zentrum f{\"u}r Informatik.
\newblock URL: \url{https://drops.dagstuhl.de/entities/document/10.4230/LIPIcs.IPEC.2017.15}, \href {https://doi.org/10.4230/LIPIcs.IPEC.2017.15} {\path{doi:10.4230/LIPIcs.IPEC.2017.15}}.

\bibitem{durajalgorithmica2019}
Lech Duraj, Marvin K{\"{u}}nnemann, and Adam Polak.
\newblock Tight conditional lower bounds for longest common increasing subsequence.
\newblock {\em Algorithmica}, 81(10):3968--3992, 2019.
\newblock URL: \url{https://doi.org/10.1007/s00453-018-0485-7}, \href {https://doi.org/10.1007/S00453-018-0485-7} {\path{doi:10.1007/S00453-018-0485-7}}.

\bibitem{equiicalp2019}
Massimo Equi, Roberto Grossi, Veli M\"{a}kinen, and Alexandru~I. Tomescu.
\newblock {On the Complexity of String Matching for Graphs}.
\newblock In Christel Baier, Ioannis Chatzigiannakis, Paola Flocchini, and Stefano Leonardi, editors, {\em 46th International Colloquium on Automata, Languages, and Programming (ICALP 2019)}, volume 132 of {\em Leibniz International Proceedings in Informatics (LIPIcs)}, pages 55:1--55:15, Dagstuhl, Germany, 2019. Schloss Dagstuhl -- Leibniz-Zentrum f{\"u}r Informatik.
\newblock URL: \url{https://drops.dagstuhl.de/entities/document/10.4230/LIPIcs.ICALP.2019.55}, \href {https://doi.org/10.4230/LIPIcs.ICALP.2019.55} {\path{doi:10.4230/LIPIcs.ICALP.2019.55}}.

\bibitem{equisofsem2021}
Massimo Equi, Veli M{\"a}kinen, and Alexandru~I. Tomescu.
\newblock Graphs cannot be indexed in polynomial time for sub-quadratic time string matching, unless seth fails.
\newblock In Tom{\'a}{\v{s}} Bure{\v{s}}, Riccardo Dondi, Johann Gamper, Giovanna Guerrini, Tomasz Jurdzi{\'{n}}ski, Claus Pahl, Florian Sikora, and Prudence~W.H. Wong, editors, {\em SOFSEM 2021: Theory and Practice of Computer Science}, pages 608--622, Cham, 2021. Springer International Publishing.

\bibitem{equitcs2023}
Massimo Equi, Veli M{\"{a}}kinen, and Alexandru~I. Tomescu.
\newblock Graphs cannot be indexed in polynomial time for sub-quadratic time string matching, unless {SETH} fails.
\newblock {\em Theor. Comput. Sci.}, 975:114128, 2023.
\newblock URL: \url{https://doi.org/10.1016/j.tcs.2023.114128}, \href {https://doi.org/10.1016/J.TCS.2023.114128} {\path{doi:10.1016/J.TCS.2023.114128}}.

\bibitem{equitalg2023}
Massimo Equi, Veli M\"{a}kinen, Alexandru~I. Tomescu, and Roberto Grossi.
\newblock On the complexity of string matching for graphs.
\newblock {\em ACM Trans. Algorithms}, 19(3), apr 2023.
\newblock \href {https://doi.org/10.1145/3588334} {\path{doi:10.1145/3588334}}.

\bibitem{ferraginafocs2000}
P.~Ferragina and G.~Manzini.
\newblock Opportunistic data structures with applications.
\newblock In {\em Proceedings 41st Annual Symposium on Foundations of Computer Science}, pages 390--398, 2000.
\newblock \href {https://doi.org/10.1109/SFCS.2000.892127} {\path{doi:10.1109/SFCS.2000.892127}}.

\bibitem{ferraginajacm2005}
Paolo Ferragina and Giovanni Manzini.
\newblock Indexing compressed text.
\newblock {\em J. ACM}, 52(4):552–581, jul 2005.
\newblock \href {https://doi.org/10.1145/1082036.1082039} {\path{doi:10.1145/1082036.1082039}}.

\bibitem{fischeripl2010}
Johannes Fischer.
\newblock Wee lcp.
\newblock {\em Information Processing Letters}, 110(8):317--320, 2010.
\newblock URL: \url{https://www.sciencedirect.com/science/article/pii/S002001901000044X}, \href {https://doi.org/10.1016/j.ipl.2010.02.010} {\path{doi:10.1016/j.ipl.2010.02.010}}.

\bibitem{fischercpm2008}
Johannes Fischer, Veli M{\"a}kinen, and Gonzalo Navarro.
\newblock An(other) entropy-bounded compressed suffix tree.
\newblock In Paolo Ferragina and Gad~M. Landau, editors, {\em Combinatorial Pattern Matching}, pages 152--165, Berlin, Heidelberg, 2008. Springer Berlin Heidelberg.

\bibitem{fischertcs2009}
Johannes Fischer, Veli Mäkinen, and Gonzalo Navarro.
\newblock Faster entropy-bounded compressed suffix trees.
\newblock {\em Theoretical Computer Science}, 410(51):5354--5364, 2009.
\newblock Combinatorial Pattern Matching.
\newblock URL: \url{https://www.sciencedirect.com/science/article/pii/S0304397509006392}, \href {https://doi.org/10.1016/j.tcs.2009.09.012} {\path{doi:10.1016/j.tcs.2009.09.012}}.

\bibitem{fischermit1974}
M.~J. Fischer and M.~S. Paterson.
\newblock String-matching and other products.
\newblock Technical report, USA, 1974.

\bibitem{gagiesoda2018}
Travis Gagie, Gonzalo Navarro, and Nicola Prezza.
\newblock {\em Optimal-Time Text Indexing in BWT-runs Bounded Space}, pages 1459--1477.
\newblock URL: \url{https://epubs.siam.org/doi/abs/10.1137/1.9781611975031.96}, \href {https://arxiv.org/abs/https://epubs.siam.org/doi/pdf/10.1137/1.9781611975031.96} {\path{arXiv:https://epubs.siam.org/doi/pdf/10.1137/1.9781611975031.96}}, \href {https://doi.org/10.1137/1.9781611975031.96} {\path{doi:10.1137/1.9781611975031.96}}.

\bibitem{gagie2020jacm}
Travis Gagie, Gonzalo Navarro, and Nicola Prezza.
\newblock Fully functional suffix trees and optimal text searching in bwt-runs bounded space.
\newblock {\em J. ACM}, 67(1), jan 2020.
\newblock \href {https://doi.org/10.1145/3375890} {\path{doi:10.1145/3375890}}.

\bibitem{gibneyspire2020}
Daniel Gibney.
\newblock An efficient elastic-degenerate text index? not likely.
\newblock In Christina Boucher and Sharma~V. Thankachan, editors, {\em String Processing and Information Retrieval}, pages 76--88, Cham, 2020. Springer International Publishing.

\bibitem{gogexp2013}
Simon Gog and Enno Ohlebusch.
\newblock Compressed suffix trees: Efficient computation and storage of lcp-values.
\newblock {\em ACM J. Exp. Algorithmics}, 18, may 2013.
\newblock \href {https://doi.org/10.1145/2444016.2461327} {\path{doi:10.1145/2444016.2461327}}.

\bibitem{grossistoc2000}
Roberto Grossi and Jeffrey~Scott Vitter.
\newblock Compressed suffix arrays and suffix trees with applications to text indexing and string matching (extended abstract).
\newblock In {\em Proceedings of the Thirty-Second Annual ACM Symposium on Theory of Computing}, STOC '00, page 397–406, New York, NY, USA, 2000. Association for Computing Machinery.
\newblock \href {https://doi.org/10.1145/335305.335351} {\path{doi:10.1145/335305.335351}}.

\bibitem{grossiscicomp2005}
Roberto Grossi and Jeffrey~Scott Vitter.
\newblock Compressed suffix arrays and suffix trees with applications to text indexing and string matching.
\newblock {\em SIAM Journal on Computing}, 35(2):378--407, 2005.
\newblock \href {https://arxiv.org/abs/https://doi.org/10.1137/S0097539702402354} {\path{arXiv:https://doi.org/10.1137/S0097539702402354}}, \href {https://doi.org/10.1137/S0097539702402354} {\path{doi:10.1137/S0097539702402354}}.

\bibitem{gusfieldbook1997}
Dan Gusfield.
\newblock {\em Algorithms on Strings, Trees, and Sequences: Computer Science and Computational Biology}.
\newblock Cambridge University Press, 1997.

\bibitem{hoffmanjacm1982}
Christoph~M. Hoffmann and Michael~J. O'Donnell.
\newblock Pattern matching in trees.
\newblock {\em J. {ACM}}, 29(1):68--95, 1982.
\newblock \href {https://doi.org/10.1145/322290.322295} {\path{doi:10.1145/322290.322295}}.

\bibitem{indykfocs1997}
P.~Indyk.
\newblock Deterministic superimposed coding with applications to pattern matching.
\newblock In {\em Proceedings 38th Annual Symposium on Foundations of Computer Science}, pages 127--136, 1997.
\newblock \href {https://doi.org/10.1109/SFCS.1997.646101} {\path{doi:10.1109/SFCS.1997.646101}}.

\bibitem{indykfocs1998}
P.~Indyk.
\newblock Faster algorithms for string matching problems: matching the convolution bound.
\newblock In {\em Proceedings 39th Annual Symposium on Foundations of Computer Science (Cat. No.98CB36280)}, pages 166--173, 1998.
\newblock \href {https://doi.org/10.1109/SFCS.1998.743440} {\path{doi:10.1109/SFCS.1998.743440}}.

\bibitem{kalaisoda2002}
Adam Kalai.
\newblock Efficient pattern-matching with don't cares.
\newblock In {\em Proceedings of the Thirteenth Annual ACM-SIAM Symposium on Discrete Algorithms}, SODA '02, page 655–656, USA, 2002. Society for Industrial and Applied Mathematics.

\bibitem{knuthscicomp1977}
Donald~E. Knuth, James~H. Morris, Jr., and Vaughan~R. Pratt.
\newblock Fast pattern matching in strings.
\newblock {\em SIAM Journal on Computing}, 6(2):323--350, 1977.
\newblock \href {https://arxiv.org/abs/https://doi.org/10.1137/0206024} {\path{arXiv:https://doi.org/10.1137/0206024}}, \href {https://doi.org/10.1137/0206024} {\path{doi:10.1137/0206024}}.

\bibitem{kociumakaalgorithmica2019}
Tomasz Kociumaka, Jakub Radoszewski, and Tatiana Starikovskaya.
\newblock Longest common substring with approximately k mismatches.
\newblock {\em Algorithmica}, 81(6):2633--2652, 2019.
\newblock URL: \url{https://doi.org/10.1007/s00453-019-00548-x}, \href {https://doi.org/10.1007/S00453-019-00548-X} {\path{doi:10.1007/S00453-019-00548-X}}.

\bibitem{kosarajufocs1989}
S.R. Kosaraju.
\newblock Efficient tree pattern matching.
\newblock In {\em 30th Annual Symposium on Foundations of Computer Science}, pages 178--183, 1989.
\newblock \href {https://doi.org/10.1109/SFCS.1989.63475} {\path{doi:10.1109/SFCS.1989.63475}}.

\bibitem{makinenbook2023}
Veli M{\"{a}}kinen, Djamal Belazzougui, Fabio Cunial, and Alexandru~I. Tomescu.
\newblock {\em Genome-Scale Algorithm Design: Bioinformatics in the Era of High-Throughput Sequencing (2nd edition)}.
\newblock Cambridge University Press, 2023.
\newblock URL: \url{http://www.genome-scale.info/}.

\bibitem{manbersoda1990}
Udi Manber and Gene Myers.
\newblock Suffix arrays: a new method for on-line string searches.
\newblock In {\em Proceedings of the First Annual ACM-SIAM Symposium on Discrete Algorithms}, SODA '90, page 319–327, USA, 1990. Society for Industrial and Applied Mathematics.

\bibitem{manberscicomp1993}
Udi Manber and Gene Myers.
\newblock Suffix arrays: A new method for on-line string searches.
\newblock {\em SIAM Journal on Computing}, 22(5):935--948, 1993.
\newblock \href {https://arxiv.org/abs/https://doi.org/10.1137/0222058} {\path{arXiv:https://doi.org/10.1137/0222058}}, \href {https://doi.org/10.1137/0222058} {\path{doi:10.1137/0222058}}.

\bibitem{manber1992}
Udi Manber and Sun Wu.
\newblock Approximate string matching with arbitrary costs for text and hypertext.
\newblock In {\em Advances In Structural And Syntactic Pattern Recognition}, pages 22--33. World Scientific, 1992.

\bibitem{munrojalg2001}
J.Ian Munro, Venkatesh Raman, and S.Srinivasa Rao.
\newblock Space efficient suffix trees.
\newblock {\em Journal of Algorithms}, 39(2):205--222, 2001.
\newblock URL: \url{https://www.sciencedirect.com/science/article/pii/S0196677400911519}, \href {https://doi.org/10.1006/jagm.2000.1151} {\path{doi:10.1006/jagm.2000.1151}}.

\bibitem{navarro2000}
Gonzalo Navarro.
\newblock Improved approximate pattern matching on hypertext.
\newblock {\em Theor. Comput. Sci.}, 237(1–2):455–463, apr 2000.
\newblock \href {https://doi.org/10.1016/S0304-3975(99)00333-3} {\path{doi:10.1016/S0304-3975(99)00333-3}}.

\bibitem{navarrobook2016}
Gonzalo Navarro.
\newblock {\em Compact Data Structures: A Practical Approach}.
\newblock Cambridge University Press, 2016.

\bibitem{navarrosurveys2007}
Gonzalo Navarro and Veli M\"{a}kinen.
\newblock Compressed full-text indexes.
\newblock {\em ACM Comput. Surv.}, 39(1):2–es, apr 2007.
\newblock \href {https://doi.org/10.1145/1216370.1216372} {\path{doi:10.1145/1216370.1216372}}.

\bibitem{navarrodcc2014}
Gonzalo Navarro and Luís M.~S. Russo.
\newblock Fast fully-compressed suffix trees.
\newblock In {\em 2014 Data Compression Conference}, pages 283--291, 2014.
\newblock \href {https://doi.org/10.1109/DCC.2014.40} {\path{doi:10.1109/DCC.2014.40}}.

\bibitem{ohlebuschspire2010}
Enno Ohlebusch, Simon Gog, and Adrian K{\"u}gel.
\newblock Computing matching statistics and maximal exact matches on compressed full-text indexes.
\newblock In Edgar Chavez and Stefano Lonardi, editors, {\em String Processing and Information Retrieval}, pages 347--358, Berlin, Heidelberg, 2010. Springer Berlin Heidelberg.

\bibitem{park1995}
Kunsoo Park and Dong~Kyue Kim.
\newblock String matching in hypertext.
\newblock In Zvi Galil and Esko Ukkonen, editors, {\em Combinatorial Pattern Matching}, pages 318--329, Berlin, Heidelberg, 1995. Springer Berlin Heidelberg.

\bibitem{patrascustoc2010}
Mihai Patrascu.
\newblock Towards polynomial lower bounds for dynamic problems.
\newblock In {\em Proceedings of the Forty-Second ACM Symposium on Theory of Computing}, STOC '10, page 603–610, New York, NY, USA, 2010. Association for Computing Machinery.
\newblock \href {https://doi.org/10.1145/1806689.1806772} {\path{doi:10.1145/1806689.1806772}}.

\bibitem{polakipl2018}
Adam Polak.
\newblock Why is it hard to beat o(n2) for longest common weakly increasing subsequence?
\newblock {\em Information Processing Letters}, 132:1--5, 2018.
\newblock URL: \url{https://www.sciencedirect.com/science/article/pii/S0020019017302016}, \href {https://doi.org/10.1016/j.ipl.2017.11.007} {\path{doi:10.1016/j.ipl.2017.11.007}}.

\bibitem{rautiainen2017}
Mikko Rautiainen and Tobias Marschall.
\newblock Aligning sequences to general graphs in o(v + me) time.
\newblock {\em bioRxiv}, 2017.
\newblock URL: \url{https://www.biorxiv.org/content/early/2017/11/08/216127}, \href {https://arxiv.org/abs/https://www.biorxiv.org/content/early/2017/11/08/216127.full.pdf} {\path{arXiv:https://www.biorxiv.org/content/early/2017/11/08/216127.full.pdf}}, \href {https://doi.org/10.1101/216127} {\path{doi:10.1101/216127}}.

\bibitem{russolata2008}
Lu{\'i}s M.~S. Russo, Gonzalo Navarro, and Arlindo~L. Oliveira.
\newblock Fully-compressed suffix trees.
\newblock In Eduardo~Sany Laber, Claudson Bornstein, Loana~Tito Nogueira, and Luerbio Faria, editors, {\em LATIN 2008: Theoretical Informatics}, pages 362--373, Berlin, Heidelberg, 2008. Springer Berlin Heidelberg.

\bibitem{russotalg2011}
Lu\'{\i}s M.~S. Russo, Gonzalo Navarro, and Arlindo~L. Oliveira.
\newblock Fully compressed suffix trees.
\newblock {\em ACM Trans. Algorithms}, 7(4), sep 2011.
\newblock \href {https://doi.org/10.1145/2000807.2000821} {\path{doi:10.1145/2000807.2000821}}.

\bibitem{sadakanesoda2002}
Kunihiko Sadakane.
\newblock Succinct representations of lcp information and improvements in the compressed suffix arrays.
\newblock In {\em Proceedings of the Thirteenth Annual ACM-SIAM Symposium on Discrete Algorithms}, SODA '02, page 225–232, USA, 2002. Society for Industrial and Applied Mathematics.

\bibitem{sadakanetcsyst2007}
Kunihiko Sadakane.
\newblock Compressed suffix trees with full functionality.
\newblock {\em Theory Comput. Syst.}, 41(4):589--607, 2007.
\newblock URL: \url{https://doi.org/10.1007/s00224-006-1198-x}, \href {https://doi.org/10.1007/S00224-006-1198-X} {\path{doi:10.1007/S00224-006-1198-X}}.

\end{thebibliography}

\documentclass[envcountsect,11pt,letterpaper]{llncs}
\usepackage{soul}
\usepackage[utf8]{inputenc}
\usepackage[table]{xcolor}
 \usepackage{cancel} 
\usepackage{float}
\usepackage[margin=1in]{geometry}
\usepackage{graphicx}
\usepackage{xcolor}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage[sort,nocompress]{cite}
\usepackage[ruled,vlined,lined,boxed,commentsnumbered]{algorithm2e}
\usepackage{todonotes}
\usepackage{verbatim}
\usepackage{subcaption}
\usepackage[normalem]{ulem}
\usepackage{array}
\usepackage{algorithmic}
\usepackage{mathrsfs}
 \usepackage{multirow}
 \usepackage{blindtext}
\usepackage{hyperref}
\usepackage{graphicx}
\usepackage{todonotes}

\usepackage{pgf}
\usepackage{tikz}
\usepackage{tikz-cd}
\tikzcdset{scale cd/.style={every label/.append style={scale=#1},
    cells={nodes={scale=#1}}}}

\usetikzlibrary{arrows,automata,positioning}
\SetKwComment{Comment}{$\triangleright$\ }{}

\newcommand{\LCP}{\mathsf{LCP}}
\newcommand{\lcp}{\mathop{\mathsf{lcp}}}

\title{Enhanced Graph Pattern Matching}

\author{
Nicola Cotumaccio
} 

\institute{
}

\date{\today}

\begin{document}

\maketitle
\thispagestyle{empty}

\begin{abstract}
Pattern matching queries on strings can be solved in linear time by Knuth–Morris–Pratt (KMP) algorithm. In 1973, Weiner introduced the suffix tree of a string [FOCS 1973] and showed that the seemingly more difficult problem of computing matching statistics can also be solved in liner time. Pattern matching queries on graphs are inherently more difficult: under the Orthogonal Vector hypothesis, the graph pattern matching problem cannot be solved in subquadratic time [TALG 2023]. The complexity of graph pattern matching can be parameterized by the topological complexity of the considered graph, which is captured by a parameter $ p $ [JACM 2023].

In this paper, we show that, as in the string setting, computing matching statistics on graph is as difficult as solving standard pattern matching queries. To this end, we introduce a notion of longest common prefix (LCP) array for arbitrary graphs.
\end{abstract}

\setcounter{page}{1}

\section{Introduction}

The increasing amount of data has motivated a considerable effort to improve the available string processing algorithms (and in particular, the available string searching algorithm). Pattern matching is one of the fundamental problems to solve on a text, and it can be classically solved in linear time by using Knuth–Morris–Pratt (KMP) algorithm \cite{knuthscicomp1977}. Knuth-Morris-Pratt algorithm was only the start point of a terrific investigation of the problem. We can identify the following lines of research.
\begin{itemize}
    \item \emph{Enhanced pattern matching}. From the beginning of the seventies mark, there was a growing effort to understand whether it is possible to efficiently solve more complicated variants of the pattern matching problem, contributing to the development of a variety of elegant algorithmic techniques (see \cite{crochemorebook2021} for a recent book on the topic). The breakthrough result occurred in 1973, when Weiner invented \emph{suffix trees}, showing that the suffix trees support a linear-time solution of the problem (today known as the problem) of computing matching statistics. Over the years, suffix trees proved able to solve a variety of problems, including computing longest common extensions, maximal repeats, minimal absent words, maximal palindromes and the LZ77 factorization of a string (see \cite{gusfieldbook1997, makinenbook2023, navarrobook2016, apostolicocaw1985}). Some variants of suffix trees support more difficult problems, such as parameterized pattern matching \cite{bakerstoc1993}. Nonetheless, suffix trees are not suitable to handle some other natural variations, such as the string matching with don't care problem, where both the pattern and the text may contain wildcards. The classical solution of Fischer and Paterson, based on fast Fourier transform, runs in time $ O(n \log w \log \sigma) $, where $ n $ is the length of the text, $ w $ is the length of the pattern, and $ \sigma $ is the size of alphabet \cite{fischermit1974}. Indyk removed the dependence on the alphabet by giving a randomized algorithm $ O(n \log n) $ algorithm \cite{indykfocs1998}; then, Kalai gave a randomized $ O(n \log w) $ algorithm \cite{kalaisoda2002}. Finally, Cole and Hariharan gave a deterministic $ O(n \log w) $ algorithm \cite{colestoc2002, colescicomp2003}, later simplified in \cite{cliffordipl2007}.

    \item \emph{Graph pattern matching} The classic formulation of the pattern matching problem can be seen as a special instance of the problem of determining whether a string can be read by following edges in a node-labeled or edge-labeled graph, which in turn is strictly connected to the problem of deciding whether a string is accepted by a finite automaton --- in the string setting, the text string can be interpreted as a graph only consisting of a single path. The graph pattern matching problem was initially motivated by the need of retrieve information stored in hypertext \cite{manber1992, tasuya1993, park1995, amir2000, rautiainen2017, navarro2000}. Assuming a constant alphabet, pattern matching on labeled trees can be solved in linear time \cite{tasuya1993}. Amir et al. considered the general setting in which nodes can be labeled with nonempty strings, showing that pattern matching queries on arbitrary graphs can be solved in $ O(\mathfrak{e} + |\pi| e) $ time, where $ e $ is the number of edges in the graph, $ |\pi| $ is the length of the pattern, and $ \mathfrak{e} $ is the total length of all labels in the graph \cite{amir2000}.

    A related extension is the \emph{tree pattern matching problem}. In the literature, the three pattern matching problem is not the problem of determining whether a string occurs in a tree, but it is the problem of deciding whether a labeled tree of size $ w $ occurs as a subtree of an another labeled tree of size $ n $. The tree pattern matching problem finds a variety of applications, including symbolic computation, code optimization of compilers and implementations of abstract data types \cite{hoffmanjacm1982}. The naive algorithm runs in $ O(wn) $ time. Kosaraju designed an $ \tilde{O}(w^{0.75}n) $ algorithm \cite{kosarajufocs1989}; Dubiner et al. gave an $ \tilde{O}(\sqrt{w} n) $ algorithm \cite{dubinerjacm1994, dubinerfocs1990}; Cole and Hariharan exhibited an $ O(n \log w) $ algorithm through a reduction to the subset matching problem \cite{colestoc2002, colescicomp2003} (see also \cite{indykfocs1997, indykfocs1998}). The tree matching problem stimulated many new algorithmic techniques; for example, in his paper, Kosaraju introduced the suffix tree of a tree and the convolution of a tree and a string \cite{kosarajufocs1989}.
    \item \emph{Compressed pattern matching} The field of compressed data structures aims to compress data in such a way that it is possible to retrieve information without needing to decompress the stored data \cite{navarrobook2016, makinenbook2023, adjerohbook2008}. Being able to solve pattern matching on \emph{compressed} data is arguably of the most important queries, and possibly the most studied problem in the compressed data realm. Storing the suffix tree of a string requires $ O(n \log n) $ bits (where $ n $ is length of the string), and the hidden constant in the space bound is generally too big for big data applications. Manber and Myers introduced \emph{suffix arrays}, which do not have the full functionality of suffix trees, but only require $ n \log n $ bits while still supporting conventional pattern matching queries \cite{manberscicomp1993, manbersoda1990}. The year 2000 was a crucial one for compressed pattern matching: on the one hand, Grossi and Vitter showed how to compress the text and the suffix array while supporting string searching queries \cite{grossistoc2000, grossiscicomp2005}; on the other hand, Ferragina and Manzini invented the FM-index \cite{ferraginafocs2000, ferraginajacm2005}, which also supports string searching queries on a compressed representation based on the Burrows-Wheeler Transform \cite{burrows1994}. Subsequently, Sadakane showed how to compress \emph{suffix trees} while retaining their full functionality \cite{sadakanetcsyst2007, sadakanesoda2002}, notably improving on previous work \cite{munrojalg2001, grossiscicomp2005}. Subsequently, a plethora of (slightly different) compressed suffix trees were proposed, with different theoretical and practical space-time tradeoffs (for example, \cite{fischercpm2008, fischertcs2009, fischeripl2010, russotalg2011, russolata2008, gogexp2013, abeliukalgorithms2013, navarrodcc2014}), which also depend on the specific variant of compressed suffix array used in the compressed suffix tree (see \cite{navarrosurveys2007}). Recently, Gagie et al. showed how to design a fully-functional suffix tree with space and time bound parameterized by the number $ r $ of runs in the Burrows-Wheeler transform \cite{gagie2020jacm, gagiesoda2018}.
\end{itemize}

Distinct variants of the pattern matching problem have distinct complexities and require notably different approaches, which suggests that these problems can represent a benchmark to test the limits of algorithmic techniques. When it comes to lower bounds, a recent and fruitful line of research has shown that the complexity of string processing problems can be often determined assuming (some variant of) the Strong Exponential Time Hypothesis or the Orthogonal Vector Hypothesis. Under these conjectures, multiple problems cannot be solved in strongly subquadratic time, including computing local alignments \cite{abboudicalp2014}, the Fréchet distance \cite{bringmannfocs2014}, the edit distance \cite{backursscicomp2018, backursstoc2015}, the dynamic time warping distance and longest common subsequences \cite{abboudfocs2015}, longest palindromic subsequences and longest tandem subsequences \cite{bringmannfocs2015}. These techniques were adapted to prove the conditional hardness of several other problems, including longest common increasing subsequences \cite{durajalgorithmica2019, durajipec2017}, elastic-degenerate text \cite{gibneyspire2020}, episode matching \cite{billecpm2022}, Wheeler languages \cite{beckerspire2023} (see \cite{patrascustoc2010, amirfocs2014, abboudstocs2016, backursfocs2016, polakipl2018, kociumakaalgorithmica2019} for more conditional results with the same flavor).

When it comes to lower bounds for pattern matching problems (conditional to the Strong Exponential Time Hypothesis or the Orthogonal Vector Hypothesis), once again small variation of the problem can lead to drastic differences. For example, while the tree pattern matching problem can be solved in quasilinear time (as we saw before), if we consider the unlabeled case --- that is, if we consider the subtree isomorphism problem --- then the problem cannot be solved in subquadratic time \cite{abboudtalg2018, abboudsoda2016}. Recently, Equi et al. showed that the graph pattern matching problem cannot be solved in subquadratic time \cite{equitalg2023, equiicalp2019, equitcs2023, equisofsem2021}.

\section{Our results}

In view of Equi et al. result, the pattern matching problem on graphs is inherently more difficult that the pattern matching on strings. Cotumaccio et al. introduced a parameter $ p $ associated to a graph that refines the complexity of pattern matching --- between linear and quadratic, depending on $ p $ \cite{cotumacciojacm2023, cotumacciosoda2021}. The natural question is whether more complex variant problem of graph pattern matching problem are still tractable or not.

Let us consider the problem originally considered by Weiner when introducing suffix trees: computing matching statistics. Weiner showed that, even though the problem of computing matching statistics appears more difficult that conventional pattern matching, it can still be solved in linear time by using a suffix tree.

We can prove that, even on graphs, matching statistics can be computed as efficiently as solving conventional pattern matching queries. Here we will only provide the main intuition needed to prove our claim; the full proof is quite technical, and will be provided in the extended version of this paper.

In the string setting, matching statistics can be computed using the Burrows-Wheeler Transform and the LCP array \cite{ohlebuschspire2010}. Since Cotumaccio et al.'s work generalizes the Burrows-Wheeler Transform to graphs, our first step is to generalize the LCP array to graphs. Consider a graph $ G $ with $ n $ nodes, and assume without loss of generality that  every node of $ G $ has at least one incoming edge. For every node $ u $, we consider the lexicographically smallest string $ \min_u $ in $ \Sigma^\omega $ that can be read starting from $ u $ and following edges in a backward fashion. We sort all the strings $ \min_u $'s, obtain strings $ \alpha_1 \preceq \alpha_2 \preceq \dots \preceq \alpha_n $. Then, we define the array $ \LCP^{\min}_G[2, n] $ such that $ \LCP^{\min}_G[i] = \lcp(\alpha_{i - 1}, \alpha_i) $. Analogously, we define an array $ \LCP^{\max}_G[2, n] $ by considering the lexicographically \emph{largest} string reaching each node. Note that it would be possible to store more information about $ G $ by defining a \emph{unique} LCP array of size $ 2n - 1 $ obtained by jointly sorting the minimum and the maximum string reaching each node.
 
In the graph setting, however, the LCP array is not sufficient, because each node is potentially reached by infinitely many paths. The crux of our proof is to show that it is possible to mantain $ O(p) $ values during the execution of our algorithm in such a way that (i) we can compute matching statistics and (ii) the $ O(p) $ values can be quickly updated.

In the full version we will prove the following theorem.

\begin{theorem}
    Let $ G $ be a graph with parameter $ 1 \le p \le n $ \cite{cotumacciojacm2023}, where $ n $ is the number of vertices. Then, there exists a data structure such that, given a string of length $ w $, we can compute the matching statistics of $ w $ with respect to $ G $ in $ O(wp^2 \log \log (p \sigma)) $ time, where $ \sigma $ is the size of the alphabet.
\end{theorem}

\bibliographystyle{plainurl}
\bibliography{main.bib}

\end{document}